A mortgage of $400,000 is to be repaid over a 10 year period at fixed interest rate of 5%. Determine the monthly payments.
Answer #1

Loan amount = $ 400,000

Time = 10 years = 10 × 12 = 120 months

Interest rate = 5% per annum

Let us assume the monthly payment is $ A.

\large A = 400,000(A/P, \frac{5}{12}\%,120)\\ \\ \implies A = 400,000\times \frac{0.05/12}{1-(1+0.05/12)^{-120}}\\ \\ \implies A = 400,000 \times 0.0106065\\ \\ \implies A \approx \$ \ 4,242.62

Monthly payment = $ 4,242.62.

Please contact if having any query will be obloged to you for your generous support. Your help mean a lot to me, please help. Thank you.